OpenStudy (anonymous):

A sample of Geometry class final exam grades are as follows: 80, 85, 72, 76, 95 Find the variance.

OpenStudy (anonymous):

what is variance

OpenStudy (amistre64):

variance is the standarad deviation without the square root :)

OpenStudy (anonymous):

o

OpenStudy (amistre64):

mean = 81.6 variance =(8.90505)^2

OpenStudy (amistre64):

80-81.6 = .... alot of hard work lol 85-81.6 = 72-81.6 = 76-81.6 = 95-81.6 =

OpenStudy (anonymous):

so subtract them

OpenStudy (amistre64):

its about 79.3

OpenStudy (anonymous):

rounded or not

OpenStudy (amistre64):

subtract them; square them; add them; divide by n-1 where n is the number of data points collected; and there ya go :)

OpenStudy (amistre64):

rounded; but i dont know how your answer looks, you got options?

OpenStudy (anonymous):

i think im gonna give up on this problem

OpenStudy (anonymous):

1) 63.44 2) 7.96 3) 49.23 4) 6.72

OpenStudy (amistre64):

yeah, one down aint goona hurt much ;)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

and those are the answers i have

OpenStudy (anonymous):

i know how you feel these stats problems are a lot of tedious work

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

some calculators make it a bit easier

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

if we are doing population variance; it closer to 64

OpenStudy (amistre64):

choose the first option :)

OpenStudy (anonymous):

i jst guessed